# NEZHA: A Zero-sacrifice and Hyperspeed Decoding Architecture for Generative Recommendations

### v1

@@ 3.4. Training and Inference
This section details the training and inference procedures for NEZHA, which are summarized in Algorithm 3. While both processes share the same core architecture, their objectives differ, leading to distinct approaches for token selection and state transition.
During training, the model learns to predict the ground-truth semantic ID using a teacher-forcing approach. For a given training instance with label $y=[t^{y}_{1},\dots,t^{y}_{L}]$: LLM first initializes the representations with a single call(line 2). At each step $l$, the model uses Equation (2) to compute the probability distribution $\bm{p}_{l}$ (line 5). For the subsequent transition step (Equation (3)), the context $\bm{s}_{l}$ is updated using the embedding of the ground-truth token, i.e., $\bm{e}_{l}=\bm{e}^{y}_{l}$ (line 6-7). This ensures the model learns the correct sequential path, regardless of its own predictions at step $l$. Finally, the probability corresponding to the ground-truth token, $p^{y}_{l}$, is used to compute the cross-entropy loss, which is minimized across the entire dataset $\mathcal{D}$ (line 9-10):
@@
In contrast, the goal for the inference is to find the most probable semantic IDs, for which we employ beam search. The process begins by prefilling the specialized prompt (line 13). Then, we initialize the step index $l$ and context state $\bm{s}_{1}$, set up the prediction set $\hat{Y}_{0}$ and state set $\mathcal{S}_{0}$ to track the candidate beams and their respective states (line 14). For each step, we iterate each candidate beam (line 17), computing the probability distribution $\bm{p}_{l}$ and selecting the top $K$ tokens $t^{k}_{l}$ with the highest probabilities $p^{k}_{l}$ (line 18-19). The context state for each of beam is then updated independently using the embedding of its selected token $\bm{e}^{k}_{l}$ as in line 20. New beams are finally included to $\hat{Y}$ and $\mathcal{S}_{l}$ (line 21) for further pruning (line 27), where the top $K$ overall sequences are retained at each step based on their cumulative probabilities. Notably, the final decoding step ($l=L$) incorporates a crucial verification stage to filter out hallucinated (i.e., invalid) semantic IDs, instead of simply selecting the top candidates by probability (line 23-26).
We apply the verification step only at the end because we empirically observed that hallucination rates are negligible in early steps but increase sharply at the final token. However, given the near-zero cost of our lookup-based verification, it could be applied at every step to ensure validity across the inference without a significant efficiency penalty.
@@ 4.1. Settings
Datasets. We evaluate NEZHA on three public datasets widely used as benchmarks for GR: Yelp, Amazon Beauty, and Amazon Games (Wang et al., 2024; Zheng et al., 2024). These datasets are particularly suitable as they contain rich contextual information, such as textual item descriptions. Following standard practice in prior work, we employ a leave-one-out strategy for data splitting. For each user, the last interaction is reserved for the test, the penultimate interaction is used for validation, and all remaining interactions constitute the training set.
Baselines and Backbones. We construct the generative recommender with two backbones, Llama-1B (Dubey et al., 2024) and QWen3-0.8B (Yang et al., 2025b), which have a similar size to the LLM we used in our production environments. We primarily compare with two baselines on public datasets: vanilla LLM decoding with beam search and MTP (Gloeckle et al., 2024). Other baselines, such as Medusa (Cai et al., 2024; Zagyva et al., 2025) and AtSpeed (Lin et al., 2025), are omitted due to their excessive latency and complexity for practical use.
@@ 4.3. Ablation Study
Additionally, we present the ablation results on offline production data in Figure 3 to further validate our conclusions. Notably, NEZHA consistently outperforms all variants. However, the trends differ for specific variants when compared to results from public data, which may stem from variations in data distribution and volume. For instance, the volume of items significantly increases in production data, resulting in greater performance degradation for NEZHA-3 and NEZHA-4. This degradation can be led by diminishing the sequential dependency and introducing too many invalid predictions, respectively. Meanwhile, NEZHA-1 and NEZHA-2 show the consistent pattern as on public datasets: NEZHA-1 is omitted from the figure due to its limited performance, while NEZHA-2 is only slightly worse than the original version.
Figure 4. Illustration of the method’s deployment in Taobao Search. The example shows results for the query “Dress” (top left, in the red box). Advertisements are identified by a label in the bottom-right corner of each product, in the orange box. Sensitive information has been obscured for confidentiality.
@@ 5. Real-world Deployment
The efficacy of this deployment is confirmed through extensive experiments. Offline evaluations show an absolute increase in the hit rate on clicked items by 0.58% and 0.61% for the top 500 and 1000 results, respectively. The positive result enables the deployment of the GR system in the prime position as of October 2025, where 7-day online A/B testing registered a 1.2% revenue increase for a billion-level business.
@@ 6. Related Works
Generative Recommendations. Significant research has been devoted to advancing each stage of the GR pipeline, i.e., item tokenization, LLM training, and LLM serving. For item tokenization, beyond using multi-modal content, recent work has started to incorporate collaborative information to create more effective semantic IDs (Wang et al., 2024; Zhu et al., 2024). The design of these IDs has also been extended to more complex scenarios, such as multi-domain (Zheng et al., 2025b) and multi-behavior (Liu et al., 2024a) recommendations, while other approaches seek to unify the training of the tokenizer and the LLM in an end-to-end framework (Liu et al., 2024b). For training, researchers have explored pre-training paradigms to align item semantics with the LLM’s linguistic space (Zheng et al., 2024), as well as post-training techniques like Direct Preference Optimization (DPO) (Rafailov et al., 2023) to fine-tune model outputs toward specific business objectives (Deng et al., 2025; Wei et al., 2025). To facilitate serving, researchers have adapted speculative decoding for top-k recommendation (Lin et al., 2025), enabled out-of-vocabulary inference with SpecGR (Ding et al., 2024), and reformulated the system with long semantic IDs and graph-based parallel decoding in RPG (Hou et al., 2025).
Despite their impressive performance, the practical deployment of GR models is severely hampered by high inference latency. The autoregressive generation of semantic IDs requires multiple sequential forward passes through the LLM, a process that is often too slow for industrial systems demanding real-time responses, motivating the proposal of NEZHA.
Speculative Decoding. To address the high latency associated with autoregressive generation, Speculative Decoding (SD) has emerged as a leading technique for accelerating inference (Leviathan et al., 2023; Xia et al., 2024; Zhang et al., 2025). The standard SD framework utilizes a small, fast “draft model” to generate a sequence of candidate tokens, which is then verified in a single parallel forward pass by the larger, original “target model”.
@@ 7. Conclusion
Industrial-scale Generative Recommendation (GR) using Large Language Models (LLMs) is severely constrained by high inference latency. Existing acceleration methods, including Speculative Decoding (SD), are often impractical due to their reliance on auxiliary draft models and expensive verification steps. To overcome these limitations, we propose NEZHA, a highly efficient decoding framework for GR. NEZHA introduces two key innovations: (1) self-drafting, which allows the main LLM to generate multiple candidates in a single pass, eliminating the need for a separate draft model; and (2) model-free verification, which uses a near-instantaneous hash lookup to validate candidates, bypassing costly model-based checks. Experiments show that NEZHA achieves zero-sacrifice accuracy while reducing latency to levels suitable for large-scale industrial deployment.

### v2

@@ 3.4. Training and Inference
This section details the pipeline of NEZHA, summarized in Algorithm 3, including the training and inference processes. While both processes share the architecture, their objectives differ, leading to distinct approaches for token selection and state transition.
During training, the model learns to predict the ground-truth semantic ID using a teacher-forcing approach. For a given training instance with label $y=[t^{y}_{1},\dots,t^{y}_{L}]$: LLM first initializes the representations with a single call(line 2). At each step $l$, the model uses Equation (2) to compute the probability distribution $\bm{p}_{l}$ (line 5). For the subsequent transition step (Equation (3)), the context $\bm{s}_{l}$ is updated using the embedding of the ground-truth token, i.e., $\bm{e}_{l}=\bm{e}^{y}_{l}$ (line 6-7). This ensures the model learns the correct sequential path, regardless of its own predictions at step $l$. Finally, the probability for the ground-truth token, $p^{y}_{l}$, is used to compute the cross-entropy loss, which would be minimized over $\mathcal{D}$ (line 9-10):
@@
In contrast, the goal of inference is to find the most probable semantic IDs, which we achieve using beam search. The process begins by prefilling the specialized prompt (line 13). Then, we initialize the step index $l$ and the context state $\bm{s}_{1}$, and set up the prediction set $\hat{Y}_{0}$ and the state set $\mathcal{S}_{0}$ to track the candidate beams and their respective states (line 14). For each step, we iterate over each candidate beam (line 17), compute the probability distribution $\bm{p}_{l}$, and select the top $K$ tokens $t^{k}_{l}$ with the highest probabilities $p^{k}_{l}$ (lines 18-19). The context state for each beam is then updated independently using the embedding of its selected token $\bm{e}^{k}_{l}$ as in line 20. New beams are finally included to $\hat{Y}$ and $\mathcal{S}_{l}$ (line 21) for further pruning (line 27), where the top $K$ overall sequences are retained at each step based on their cumulative probabilities. Notably, the final decoding step ($l=L$) includes a crucial verification stage to filter out hallucinated (i.e., invalid) semantic IDs, rather than simply selecting the top candidates by probability (lines 23-26).
@@ 4.1. Settings
Datasets. We evaluate NEZHA on three public datasets widely used as benchmarks for GR: Yelp, Amazon Beauty, and Amazon Games (Wang et al., 2024; Zheng et al., 2024). These datasets are particularly suitable as they contain rich contextual information, such as textual item descriptions. Following standard practice in prior work, we employ a leave-one-out strategy for data splitting. For each user, the last interaction is reserved for the test, the penultimate interaction is used for validation, and remaining interactions constitute the training set.
Baselines and Backbones. We construct the generative recommender with two backbones, Llama-1B (Dubey et al., 2024) and QWen3-0.8B (Yang et al., 2025a), which have a similar size to the LLM we used in our production environments. We primarily compare with two baselines on public datasets: vanilla LLM decoding with beam search and MTP (Gloeckle et al., 2024). Other baselines, such as Medusa (Cai et al., 2024; Zagyva et al., 2025) and AtSpeed (Lin et al., 2025), are omitted due to their excessive latency and complexity for practical use.
@@ 4.3. Ablation Study
Additionally, we present the ablation results on offline production data in Figure 2 to further validate our conclusions. Notably, NEZHA consistently outperforms all variants. However, trends differ across specific variants compared with public data, which may stem from variations in data distribution and volume. For instance, the volume of items significantly increases in production data, resulting in greater performance degradation for NEZHA-3 and NEZHA-4. This degradation can be driven by weakening the sequential dependency and by introducing too many invalid predictions, respectively. Meanwhile, NEZHA-1 and NEZHA-2 show the consistent pattern as on public datasets: NEZHA-1 is omitted from the figure due to its limited performance, while NEZHA-2 is only slightly worse than the original version.
@@ 5. Real-world Deployment
The efficacy of this deployment is confirmed through extensive experiments. Offline evaluations show an absolute increase in the hit rate on clicked items by 0.58% and 0.61% for the top 500 and 1000 results, respectively. The positive result enables the deployment of the GR system in the prime position as of October 2025, where 7-day reverse online A/B testing on 10% traffic registered a 1.2% revenue increase (billion-level improvement).
Figure 3. Illustration of the method’s deployment in Taobao Search. The example shows results for the query “Dress” (top left, in the red box). Advertisements are identified by an orange label in the bottom-right corner of each product. Sensitive information has been obscured for confidentiality.
@@ 6. Related Works
Generative Recommendations. Significant research has been devoted to advancing each stage of the GR pipeline, i.e., item tokenization, LLM training, and LLM serving. For item tokenization, beyond using multi-modal content, recent work has started to incorporate collaborative information to create more effective semantic IDs (Wang et al., 2024; Zhu et al., 2024). The design of these IDs has also been extended to more complex scenarios, such as multi-domain (Zheng et al., 2025a) and multi-behavior (Liu et al., 2024b) recommendations, while other approaches seek to unify the training of the tokenizer and the LLM in an end-to-end framework (Liu et al., 2024a). For training, researchers have explored pre-training paradigms to align item semantics with the LLM’s linguistic space (Zheng et al., 2024), as well as post-training techniques like Direct Preference Optimization (DPO) (Rafailov et al., 2023) to fine-tune model outputs toward specific business objectives (Deng et al., 2025; Wei et al., 2025). To facilitate serving, researchers adapted speculative decoding for top-k recommendation (Lin et al., 2025), enabled out-of-vocabulary inference with SpecGR (Ding et al., 2024), and reformulated the system with long semantic IDs and parallel decoding in RPG (Hou et al., 2025).
Despite the impressive performance, the practical deployment of GR models is severely hampered by high inference latency. The autoregressive generation requires multiple forward passes through LLMs, a process that is often too slow for industrial systems demanding real-time responses, motivating the proposal of NEZHA.
Speculative Decoding. To address the high latency associated with autoregressive generation, Speculative Decoding (SD) has emerged as a leading technique for accelerating inference (Leviathan et al., 2023; Xia et al., 2024; Zhang et al., 2025b). The standard SD framework utilizes a small, fast “draft model” to generate a sequence of candidate tokens, which is then verified in a single parallel forward pass by the larger, original “target model”.
@@ 7. Conclusion
Industrial-scale Generative Recommendation (GR) using Large Language Models (LLMs) is severely constrained by high inference latency. Existing acceleration methods, including Speculative Decoding (SD), are often impractical due to their reliance on auxiliary draft models and expensive verification steps. To overcome these limitations, we propose NEZHA, a highly efficient decoding framework for GR. NEZHA introduces two key innovations: (1) self-drafting, which allows the main LLM to generate multiple candidates in a single pass, eliminating the need for a separate draft model; and (2) model-free verification, which uses a near-instantaneous hash lookup to validate candidates, bypassing costly model-based checks. Experiments show that NEZHA achieves zero-sacrifice accuracy while reducing latency to levels suitable for large-scale deployment.
Acknowledgements.
This research was partially supported by National Natural Science Foundation of China (No.62502404), Hong Kong Research Grants Council (Research Impact Fund No.R1015-23, Collaborative Research Fund No.C1043-24GF, General Research Fund No.11218325), Institute of Digital Medicine of City University of Hong Kong (No.9229503), and Alibaba (CCF-Alimama Tech Kangaroo Fund No. 2024002).